If the average (mass-weighted) temperature of a column of air between the ground (p=1000 hPa) and 500 hPa is 255K, then at what height would one find the 500 hPa level?

Suppose that a cloud forms between the ground and the 500 hPa pressure level. If latent heating in the cloud contributes a warming of 10K throughout the column then how much higher (in altitude) with the 500 hPa level be compared to the answer you gave above?
Answer #1

i made one mistake p=500hpa and Po=1000hpa so p/Po=.5
and 2=e^(.01359h1) h1=51m
for case 2 same
2=e^(0.1307h2)
h2=53m